OpenStudy (anonymous):

solve the equation: (x+1)^2 = 9 Enter large zero value

OpenStudy (anonymous):

Expand the LHS.\[x^2+2 x+1=9 \]Move 9 to the LHS and then simplify.\[x^2+2 x-8=0 \] Factor the LHS.\[(x-2) (x+4)=0 \]Can you now find the two values for x?